OpenStudy (babyslapmafro):

Please check my work. Find the line tangent to y=2x^4 at x=-1.

OpenStudy (babyslapmafro):

(-1,2) y'=8x^3 m=-8 y-2=-8(x+1) Equation of the line... y=-8x-6

OpenStudy (anonymous):

-> http://openstudy.com/study#/groups/Mathematics